OpenStudy (anonymous):

the vector sum of two forces is perpendicular to their vector difference.in that case,the force is a. are equal to eachother b.are equal to eachother in magnitude c. are not equal to each other in magnitude d.can not be predicted

OpenStudy (ilovecake):

ooooo I know this one

OpenStudy (ilovecake):

I think it is c or b but I am not sure

OpenStudy (anonymous):

I think it's B

OpenStudy (anonymous):

ya its b but how

OpenStudy (anonymous):

thank you ilove cake and freewilly :)

OpenStudy (anonymous):

but may i know how?

OpenStudy (ilovecake):

Okay....... Here is a link: http://mathworld.wolfram.com/VectorSum.html

OpenStudy (ilovecake):

Does that help?

OpenStudy (anonymous):

ilovecake.....no it doesnt..i want answer to this question specifically....thanx anyways

OpenStudy (aum):

Let \(A = a_1i+a_2j\) and \(B = b_1i+b_2j\). Then, \(A+B = (a_1+b_1)i+(a_2+b_2)j\) ; \(A-B = (a_1-b_1)i+(a_2-b_2)j\) "the vector sum of two forces is perpendicular to their vector difference" Therefore, their dot product must be zero. \((A+B).(A-B) = (a_1+b_1)(a_1-b_1)+(a_2+b_2)(a_2-b_2) = \\ a_1^2-b_1^2+a_2^2 - b_2^2 = (a_1^2+a_2)^2-(b_1^2+b_2^2) = 0 \\ (a_1^2+a_2^2) = (b_1^2+b_2^2) \\ \sqrt{(a_1^2+a_2^2)} = \sqrt{(b_1^2+b_2^2)} \\ |A| = |B| \)

OpenStudy (anonymous):

thanx aum